# Eco-evolutionary dynamics of adapting pathogens and host immunity

### v2

@@ B.1 Differential equation with expiring fitness effects
We now consider that two consecutive partial sweeps of initial fitnesses $s_{1}$ and $s_{2}$ overlap if the first one is not yet at frequency $r\beta_{1}$ while the second one is already at $(1-r)\beta_{2}$. For instance, taking $r=3/4$, this an overlap occurs if the first sweep is not yet at $3/4$ of its final value while the second one is already at $1/4$ of its final value. Thus, for an overlap to occur, we need the time $\tau$ between the two partial sweeps to be smaller than $T_{r}(s_{1})-T_{1-r}(s_{2})$. For sweeps happening at rate $\rho$, this has probability $1-\exp\left(-\rho(T_{r}(s_{1})-T_{1-r}(s_{2}))\right)$. Since the two sweeps have random initial fitness effects, we find that the overall probability for two consecutive sweeps to overlap is
$P(\text{overlap})=\int_{0}^{\infty}\text{d}s_{1}\text{d}s_{2}P_{s}(s_{1})P_{s}(s_{2})\left\{1-e^{-\rho\left(T_{r}(s_{1})-T_{1-r}(s_{2})\right)}\right\}.$
@@
B.2 Distribution of partial sweep size $\beta$
@@
$P(\beta<x)$ $=P(s<-\alpha\log(1-x))$
$=s_{0}^{-1}\int_{0}^{-\alpha\log(1-x)}e^{-s/s_{0}}\text{d}s$
$=1-(1-x)^{\alpha/s_{0}}.$
@@
$P(\beta)\propto(1-\beta)^{\alpha/s_{0}-1}.$
This distribution can accomodate various shape: for $\alpha/s_{0}>1$ it peaks at 00, and for $\alpha/s_{0}<1$ it peaks at $1$. We can also compute the following formular for the moments of $\beta$:
$\langle\beta\rangle=\frac{s_{0}}{s_{0}+\alpha}\;\text{and}\;\langle\beta^{2}\rangle=\frac{s_{0}}{s_{0}+\alpha}\frac{2s_{0}}{2s_{0}+\alpha}.$
@@
B.3 Random walk: monotonous trajectories
@@
Figure S 2: Probability of a strictly monotonous trajectory in the random walk of the main text, as a function of $\beta$ (fixed) and the initial value $x_{0}$. The “exact” solution is obtained by numerically computing the product in Equation 35 up to $t=100$.
@@ B.4 Coalescent
Figure S 3: Average coalescence times $\langle T_{n}\rangle$ for a partial sweep coalescent with effective population size $N_{e}$ and a Kingman coalescent with population size $N_{e}$. For simplicity, a constant $\beta$ is used: Left: a high value $\beta=0.25$; Right: a low value $\beta=0.05$. For low $\beta$, the two coalescent processes are very similar until a high $n$. They considerably differ if $\beta$ is larger. Note that for the partial sweep process, $T_{n}$ never goes below $\rho^{-1}$.

### v3

@@ B.2 Expiring fitness effects: sweep size and probability of overlap
We now consider that two consecutive partial sweeps of initial fitnesses $s_{1}$ and $s_{2}$ overlap if the first one is not yet at frequency $r\beta_{1}$ while the second one is already at $(1-r)\beta_{2}$. In the figure of the main text, we use $r=3/4$: an overlap occurs if the first sweep is not yet at $3/4$ of its final value while the second one is already at $1/4$ of its final value. Thus, for an overlap to occur, we need the time $\tau$ between the two partial sweeps to be smaller than $T_{r}(s_{1})-T_{1-r}(s_{2})$. For sweeps happening at rate $\rho$, this has probability $1-\exp\left(-\rho(T_{r}(s_{1})-T_{1-r}(s_{2}))\right)$. Since the two sweeps have random initial fitness effects, we find that the overall probability for two consecutive sweeps to overlap is
$P_{r}(\text{overlap})=\int_{0}^{\infty}\text{d}s_{1}\text{d}s_{2}P_{s}(s_{1})P_{s}(s_{2})\left\{1-e^{-\rho\left(T_{r}(s_{1})-T_{1-r}(s_{2})\right)}\right\}.$
@@
B.3 Distribution of partial sweep size $\beta$
@@
$P(\beta<x)$ $=P(s<-\nu\log(1-x))$
$=s_{0}^{-1}\int_{0}^{-\nu\log(1-x)}e^{-s/s_{0}}\text{d}s$
$=1-(1-x)^{\nu/s_{0}}.$
@@
$P(\beta)\propto(1-\beta)^{\nu/s_{0}-1}.$
This distribution can accomodate various shape: for $\nu/s_{0}>1$ it peaks at 00, and for $\nu/s_{0}<1$ it peaks at $1$. We can also compute the following formular for the moments of $\beta$:
$\langle\beta\rangle=\frac{s_{0}}{s_{0}+\nu}\;\text{and}\;\langle\beta^{2}\rangle=\frac{s_{0}}{s_{0}+\nu}\frac{2s_{0}}{2s_{0}+\nu}.$
@@
B.4 Random walk: monotonous trajectories
@@
Figure S 4: Probability of a strictly monotonous trajectory in the random walk of the main text, as a function of $\beta$ (fixed) and the initial value $x_{0}$. The “exact” solution is obtained by numerically computing the product in Equation 41 up to $t=100$.
@@ B.5 Coalescent
Figure S 5: Average coalescence times $\langle T_{n}\rangle$ for a partial sweep coalescent with effective population size $N_{e}$ and a Kingman coalescent with population size $N_{e}$. For simplicity, a constant $\beta$ is used: Left: a high value $\beta=0.25$; Right: a low value $\beta=0.05$. For low $\beta$, the two coalescent processes are very similar until a high $n$. They considerably differ if $\beta$ is larger. Note that for the partial sweep process, $T_{n}$ never goes below $\rho^{-1}$.
@@
Appendix C Supplementary figures
Figure S 7: Example of mutation frequency trajectories that are increasing up to a frequency of $0.5$ for H3N2/HA influenza and the expiring fitness model. For the latter, parameters used are $\alpha=s_{0}=0.03$ and three values of $\rho$ to illustrate different clonal interference regimes. In each case, 10 randomly selected trajectories are plotted, with blue color indicating final loss and red final fixation.